# Characterization of the directed landscape from the KPZ fixed point

### v2

@@ 8 Exclusion processes
Finally we consider general finite range asymmetric exclusion processes (AEPs). As in the introduction, we consider a jump distribution $p$ on $\mathbb{Z}$ such that $\{v:p(v)+p(-v)>0\}$ is finite and additively generates $\mathbb{Z}$, and whose expectation satisfies $\sum_{v}vp(v)=1$. In this section, $C,c>0$ denote constants that are allowed to depend on $p$, and whose values can change from line to line.
@@
As in [QS23], we let $\operatorname{SRW}_{\epsilon}\subset\operatorname{UC}$ be the set of functions $f:\mathbb{R}\to\mathbb{R}$ such that for each $x\in\mathbb{Z}$, $f|_{[x\epsilon/2,(x+1)\epsilon/2]}$ is linear with slope ±$\pm 2\epsilon^{-1/2}$ and $f(0)\in 2\epsilon^{1/2}\mathbb{Z}$. We call $\operatorname{SRW}_{\epsilon}$ the set of simple random walk paths at scale $\epsilon$. Then $\operatorname{SRW}:=\operatorname{SRW}_{1}$ is the state space for an AEP if we extend our AEP height functions to be linear on each of the pieces $[x,x+1],x\in\mathbb{Z}$. Let $A_{\epsilon}:\operatorname{SRW}\to\operatorname{SRW}_{\epsilon}$ be the natural rescaling map given by $A_{\epsilon}f(x)=\epsilon^{1/2}f(2\epsilon^{-1}x)$.
@@ 8.1 ASEPs under various couplings
As discussed in the introduction, the nearest neighbor case where $p(v)\neq 0$ only when $v\in\{-1,1\}$ is also referred to as the asymmetric simple exclusion process (ASEP). The case where $p(1)=1,p(-1)=0$ is TASEP. For ASEP, Quastel and Sarkar [QS23] proved convergence to the KPZ fixed point.
Theorem 8.1 ([QS23, Theorem 2.2(2)]).
Consider a collection of functions $f^{\epsilon}\in\operatorname{SRW}_{\epsilon},\epsilon>0$, and $f\in\operatorname{UC}$ such that $f^{\epsilon}\to f$ in $\operatorname{UC}$. Suppose that there exists a constant $A>0$ with $f^{\epsilon}(x)\leq A(|x|+1)$ for each $\epsilon>0$ and $x\in\mathbb{R}$. Consider ASEP with jump distribution $p$. Then for any fixed $t>0$, $\mathcal{K}^{\epsilon}_{0,t}f^{\epsilon}\stackrel{{\scriptstyle d}}{{\to}}\mathfrak{h}_{t}f$ in the topology of uniform convergence on compact sets.
@@
  • Basic coupling (colored ASEP). This coupling was previously described in the introduction. To each $x\in\mathbb{Z}$, we associate two Poisson clocks, with rate $p(1)$ and $p(-1)$ respectively. All Poisson clocks are jointly independent. The rate $p(1)$ and $p(-1)$ clocks associated to $x$ encode the jumps of particles from site $x$ to $x+1$ and $x-1$, respectively, simultaneously for all instances of ASEP.
@@
    For each $\ell\in\mathbb{Z}$, we associate two Poisson clocks, with rate $p(1)$ and $p(-1)$ respectively. All such Poisson clocks are jointly independent. The rate $p(1)$ (resp. $p(-1)$) clock associated to $\ell$ encodes the right (resp. left) jumps of the particle with label $\ell$, simultaneously for all instances of ASEP.
  • Hole coupling. Instead of thinking of ASEP as particles attempting to move to the right with rate $p(1)$ and to the left with rate $p(-1)$, the model can equivalently be described as each hole attempting to move to the left with rate $p(1)$ and to the right with rate $p(-1)$, independently for all the holes. In each of the instances of ASEP, we label all the holes by integers, such that if a hole is labeled $\ell$, then the next hole to its right is labeled $\ell+1$.
    For each $\ell\in\mathbb{Z}$, we have two Poisson clocks, with rate $p(1)$ and $p(-1)$ respectively. All such Poisson clocks are jointly independent. The rate $p(1)$ (resp. $p(-1)$) clock associated to hole $\ell$ encodes the left (resp. right) jumps of this hole.
These three couplings are example of more general exotic couplings.
To define a general exotic coupling, we first introduce the following labeling of particles and holes. For any ASEP configuration, we label all particles and holes by integers via the height function. More precisely, for a particle configuration with height function $h:\mathbb{Z}\to\mathbb{Z}$, if a particle is at location $x\in\mathbb{Z}$, it is given the label $-(h(x)+x)/2$; if a hole is at location $x\in\mathbb{Z}$, it is given the label $(-h(x)+x)/2$. It is straightforward to check that this labeling rule is compatible with ASEP dynamics.
@@
For any $(a,b)\in\mathbb{N}^{2}\setminus\{(0,0)\}$, we define the $(a,b)$-exotic coupling as follows. Let $\mathbb{Z}^{2}/(a,b)$ be the quotient of $\mathbb{Z}^{2}$ under addition by $(a,b)$. We associate each element $x\in\mathbb{Z}^{2}/(a,b)$ to two Poisson clocks with rate $p(1)$ and $p(-1)$, respectively. Let $Q:\mathbb{Z}^{2}\to\mathbb{Z}^{2}/(a,b)$ denote the projection map. Any ASEP under the $(a,b)$-exotic coupling is generated by these Poisson clocks as follows. For any $\ell,k\in\mathbb{Z}$, if the particle with $\ell$ sits immediate to the left (resp. right) of the hole with label $k$, then their swap is encoded by the rate $p(1)$ (resp. $p(-1)$) clock associated to $Q(\ell,k)$.
@@ 8.2 AEPs under the basic coupling
In the remainder of this section, we consider general AEPs with the basic coupling. The setup is described in the introduction and at the beginning of Section 8. In this subsection we will prove joint convergence of AEPs to the directed landscape by appealing to Theorem 5.2. One technical claim verifying approximate monotonicity for the model is left to Section 8.3.
Throughout the section, we fix a valid jump distribution $p$, and let $\mathcal{K}^{\epsilon}=\{\mathcal{K}^{\epsilon}_{s,t}:\operatorname{SRW}_{\epsilon}\to\operatorname{SRW}_{\epsilon},s\leq t\in\mathbb{R}\}$ denote the rescaled basic coupling of AEP with this jump distribution, via (42).
As with ASEPs, KPZ fixed point convergence for general AEPs was shown in [QS23]. The convergence result applies to initial conditions which are not too far from the stationary Bernoulli product measure. We let $\nu$ be the law on $\operatorname{SRW}$ of a simple symmetric random walk path $X$ with $X(0)=0$. We let $\nu_{\epsilon}$ be the pushforward of $\nu$ under the natural map $A_{\epsilon}:\operatorname{SRW}\to\operatorname{SRW}_{\epsilon}$.
@@
  • $\lim_{\epsilon\to 0}\epsilon^{1/4}\|g_{\epsilon}d\nu_{\epsilon}\|_{L^{2}(\nu_{\epsilon})}=0$.
@@
The convergence of $\mathcal{K}^{\epsilon}_{s,t}f^{\epsilon}\stackrel{{\scriptstyle d}}{{\to}}\mathfrak{h}_{t-s}f$ is Theorem 2.2(1) in [QS23]. For the joint convergence, let $(f,\mathcal{K}_{s,t}f)$ be any joint subsequential limit in distribution of $(f^{\epsilon},\mathcal{K}^{\epsilon}_{s,t}f^{\epsilon})$. Let $U$ be any continuity set for the law of $f$ such that $\mathbb{P}[f\in U]>0$. Then $\mathbb{P}[f^{\epsilon}\in U]\to\mathbb{P}[f\in U]$ as $n\to\infty$. Moreover, since $\mathbb{P}[f\in U]>0$, if $f^{\epsilon}_{U}$ has the law of $f^{\epsilon}$ given $f^{\epsilon}\in U$, then $f^{\epsilon}_{U}$ still satisfies the three bullets above and hence the conclusion of Theorem 2.2(1) in [QS23]. Therefore for any measurable $A\subset\operatorname{UC}$, we have $\mathbb{P}[(f,\mathcal{K}_{s,t}f)\in U\times A]=\mathbb{P}[(f,\mathfrak{h}_{t-s}f)\in U\times A]$, and so $(f,\mathcal{K}_{s,t}f)\stackrel{{\scriptstyle d}}{{=}}(f,\mathfrak{h}_{t-s}f)$ as desired. ∎
@@
Take (potentially random) initial conditions $f_{0},\ldots,f_{k}\in\operatorname{UC}$ and $f_{0}^{\epsilon},\ldots,f_{k}^{\epsilon}\in\operatorname{SRW}_{\epsilon}$ for each $\epsilon>0$, such that $f_{i}^{\epsilon}\stackrel{{\scriptstyle d}}{{\to}}f_{i}$ for each $i\in\llbracket 0,k\rrbracket$. Suppose that $f_{0}^{\epsilon}\sim\nu_{\epsilon}$ and for every $i\in\llbracket 1,k\rrbracket$, the initial conditions $f_{i}$ satisfy the first two bullet points of Theorem 8.6 and that the random variables
@@
Let $\{q_{1}<\cdots<q_{k}\}$ be the set of points where $h^{\prime}$ does not exist, so that $h^{\prime}=0$ off of $[q_{1},q_{k}]$ and is constant on each interval $(q_{j},q_{j+1})$. For $i\epsilon/2\notin(q_{1},q_{k}]$ we set $v(h)_{i}=e_{i}$. Now, for each $j\in\{1,\dots,k-1\}$, let $s_{j}$ be the value of $h^{\prime}$ on $(q_{j},q_{j+1})$. For $i\epsilon/2\in(q_{j},q_{j+1}]$ we let
@@
Consider $F_{\epsilon}=A_{\epsilon}^{-1}(B_{s,n}^{\epsilon}[h])\in\operatorname{SRW}$, and let $h_{\epsilon}$ be the Radon-Nikodym derivative of its law with respect to $\nu$. Equivalently, we will show that $\|h_{\epsilon}d\nu\|_{L^{2}(\nu)}$ stays bounded as $\epsilon\to 0$. By construction, there is an integer set $Q_{\epsilon}=\{q_{1,\epsilon},\dots,q_{k,\epsilon}\}$ (up to $O(\epsilon)$-errors, this set is given by $2\epsilon^{-1}\{q_{1},\dots,q_{k}\}$ with the $q_{i}$ as above) such that conditional on $F_{\epsilon}|_{Q_{\epsilon}}$, $F_{\epsilon}$ has the same law as a simple symmetric random walk path conditioned on the values on $Q_{\epsilon}$. This follows from the fact that if we condition a parameter-$p$ Bernoulli process on $\{1,\dots,n\}$ to have exactly $k$ $1$s, then under this conditioning the $k$ $1$s are uniformly distributed amongst all $k$-element subsets of $\{1,\dots,n\}$; in particular, the parameter $p$ plays no role.
@@
The final two assumptions in Theorem 5.2 require a monotonicity claim for general AEPs. This monotonicity claim will also allow us to connect the field in (45) to the evolutions in Theorem 8.7.
For this next lemma, consider two AEPs with height functions $(h^{-}_{t})_{t\geq 0}$ and $(h^{+}_{t})_{t\geq 0}$, started at time 00 in the basic coupling. Assume that $h^{-}_{0}\leq h^{+}_{0}$. For any $t,w>0$, we denote
@@
For any $m,t,w>0$, we have $\mathbb{P}[Y_{t,w}>m]<C(t+w)\exp(-cm)+C\exp(-ct)$.
@@
Suppose that $\eta^{1}_{t},\eta^{2}_{t}$ are two AEPs with the basic coupling started from initial conditions $\eta^{1}_{0},\eta^{1}_{0}$ that agree off of a set $\{-n,\dots,n\}$. Then there exists $C_{1},C_{2}>0$ such that for all $t>0$,
@@
Let $M_{t}=\sup\{m\in\mathbb{N}:(\eta^{1}_{t}(-m),\eta^{1}_{t}(m))\neq(\eta^{2}_{t}(-m),\eta^{2}_{t}(m))\}$. Then the process $(M_{t},\eta^{1}_{t},\eta^{2}_{t})$ is jointly Markov, and since the jump distribution $p$ has finite support, the process $M_{t}$ is stochastically dominated by $N_{t}+n$, where $N_{t},t\geq 0$ is a Poisson counting process with some $p$-dependent rate $C_{1}/2>0$. The bound in the lemma then follows from a standard tail estimate on Poisson random variables. ∎
@@
    $\mathcal{K}_{t_{*},t}(B_{t_{*},1}+h_{m})\geq\tilde{\mathcal{K}}_{0,t}f_{i}\qquad$ for $\text{for each }m\in\mathbb{N},t\in T_{*},t>t_{*}.$
@@
where $M_{\epsilon},\epsilon>0$ is a tight collection of random variables and $N_{\epsilon}\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}0$ with $\epsilon$. Here the convergence of $N_{\epsilon}$ to 00 uses Lemma 8.10 and the tightness of $M_{\epsilon}$ uses Lemma 8.9. The inequality (46) follows by taking $\epsilon\to 0$.
@@
is tight. If we take any subsequential limit $S$, jointly with (45), then by applying Lemmas 8.9 and 8.10 as before we would have that $\mathcal{K}_{r,t}g\leq\mathcal{K}_{s,t}f+S$. This would imply part $2$ if we could show that $S=\sup(g-\mathcal{K}_{s,r}f)$. However, this is not obvious and so we work around this difficulty with a truncation argument.
For each $m\in\mathbb{N}$, define $w_{m}\in\mathscr{H}$ by letting $w_{m}(x)=-1/m$ when $x\in[-m,m]$, and $w_{m}(x)=-m$ when $x\in[-m-1,m+1]^{c}$, and letting $w_{m}$ be linear on $[-m-1,-m]$ and $[m,m+1]$. Then similarly, the collection of random variables
@@
with probability $\to 1$ as $\epsilon\to 0$. Then by using Lemmas 8.9 and 8.10 as before we have that
@@
For part $3$, we let $R^{\epsilon}=\sup(f_{i}^{\epsilon}-f_{0}^{\epsilon})$ and note that $(R^{\epsilon})_{\epsilon>0}$ is again tight, as assumed in Theorem 8.7. Then using Lemmas 8.9 and 8.10 as in previous parts,
@@
where $E_{\epsilon}\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}0$ as $\epsilon\to 0$. Note that $\mathcal{K}_{0,t_{*}}^{\epsilon}f_{0}^{\epsilon}(0)$ is also tight according to Theorem 8.6. Now let $R$ be a joint subsequential limit of $R^{\epsilon}+\mathcal{K}_{0,t_{*}}^{\epsilon}f_{0}^{\epsilon}(0)$ taken with the other random variables. The above display together with the monotonicity argument from previous parts (again using Lemmas 8.9 and 8.10) implies that if $h\in\mathscr{H}$ is any function with $h(x)\geq R+1$ for all large enough $|x|$, and $B_{t_{*},1}(x)+h(x)>\tilde{\mathcal{K}}_{0,t_{*}}f_{i}(x)$ for all $x$, then
@@
for all $t\in T_{*}$, $t>t_{*}$. Finally, by choosing any decreasing sequence $\{h_{m}\}_{m\in\mathbb{N}}$ satisfying $B_{t_{*},1}+h_{m}\downarrow\tilde{\mathcal{K}}_{0,t_{*}}f_{i}$ yields the conclusion. (Here the existence of such a sequence is guaranteed by the fact that both $B_{t_{*},1}$ and $\mathcal{K}_{0,t_{*}}f_{i}$ are continuous, due to the fact that $\mathcal{K}_{0,t_{*}}f_{i}\stackrel{{\scriptstyle d}}{{=}}\mathfrak{h}_{t_{*}}f_{i}$ by Theorem 8.6, and that $B_{t_{*},1}$ is a Brownian motion.) ∎

### v3

@@ 7 Exclusion processes
Finally we consider general finite range asymmetric exclusion processes (AEPs). Consider a jump distribution $p$ on $\mathbb{Z}$ such that $\{v:p(v)+p(-v)>0\}$ is finite and additively generates $\mathbb{Z}$, and whose expectation satisfies $\sum_{v}vp(v)=1$.
@@
As in [QS23], we let $\operatorname{SRW}_{\epsilon}\subset\operatorname{UC}$ be the set of functions $f:\mathbb{R}\to\mathbb{R}$ such that for each $x\in\mathbb{Z}$, $f|_{[x\epsilon/2,(x+1)\epsilon/2]}$ is linear with slope $\pm 2\epsilon^{-1/2}$ and $f(0)\in 2\epsilon^{1/2}\mathbb{Z}$. We call $\operatorname{SRW}_{\epsilon}$ the set of simple random walk paths at scale $\epsilon$. Then $\operatorname{SRW}:=\operatorname{SRW}_{1}$ is the state space for an AEP if we extend our AEP height functions to be linear on each of the pieces $x,x+1],x\in\mathbb{Z}$. Let $A_{\epsilon}:\operatorname{SRW}\to\operatorname{SRW}_{\epsilon}$ be the natural rescaling map given by $A_{\epsilon}f(x)=\epsilon^{1/2}f(2\epsilon^{-1}x)$.
@@ 7.1 ASEPs under various couplings
As discussed in the introduction, the nearest neighbor case where $p(v)\neq 0$ only when $v\in\{-1,1\}$ is also referred to as the asymmetric simple exclusion process (ASEP). The case where p(1)=1,p(1)=0p(1)=1,p(-1)=0 is TASEP. For ASEP, Aggarwal, Corwin, and Hegde [ACH24b, ACH24a] proved convergence to the KPZ fixed point.
Theorem 7.1.
Consider a collection of functions $f^{\epsilon}\in\operatorname{SRW}_{\epsilon},\epsilon>0$, and $f\in\operatorname{UC},f\not\equiv$-$\infty$ such that $f^{\epsilon}\to f$ in $\operatorname{UC}$. Suppose that there exists a constant $A>0$ with $f^{\epsilon}(x)\leq A(|x|+1)$ for each $\epsilon>0$ and $x\in\mathbb{R}$. Consider ASEP with jump distribution $p$. Then for any fixed $t>0$, $\mathcal{K}^{\epsilon}_{0,t}f^{\epsilon}\stackrel{{\scriptstyle d}}{{\to}}\mathfrak{h}_{t}f$ in the topology of uniform convergence on compact sets.
In the special case of TASEP, this is the main theorem of [MQR21]. For general ASEP, this is proven as part of Theorem 2.10 in [ACH24a]. That theorem actually shows the stronger claim that when $k$ ASEP evolutions are coupled together in the basic coupling, then evolutions converge jointly to $k$ KPZ fixed points coupled through the directed landscape. One consequence of the upcoming Theorem 7.4 is a proof of this in the special case of basic coupled TASEP that bypasses the Yang-Baxter framework of [ACH24b, ACH24a].
@@
  • Basic coupling (colored ASEP). This coupling was previously described in the introduction. To each $x\in\mathbb{Z}$, we associate two Poisson clocks, with rate $p(1)$ and p(1)p(-1) respectively. All Poisson clocks are jointly independent. The rate $p(1)$ and p(1)p(-1) clocks associated to $x$ encode the jumps of particles from site $x$ to $x+1$ and x1x-1, respectively, simultaneously for all instances of ASEP.
@@
    For each $\ell\in\mathbb{Z}$, we associate two Poisson clocks, with rate $p(1)$ and p(1)p(-1) respectively. All such Poisson clocks are jointly independent. The rate $p(1)$ (resp. p(1)p(-1)) clock associated to $\ell$ encodes the right (resp. left) jumps of the particle with label $\ell$, simultaneously for all instances of ASEP.
  • Hole coupling. Instead of thinking of ASEP as particles attempting to move to the right with rate $p(1)$ and to the left with rate p(1)p(-1), the model can equivalently be described as each hole attempting to move to the left with rate $p(1)$ and to the right with rate p(1)p(-1), independently for all the holes. In each of the instances of ASEP, we label all the holes by integers, such that if a hole is labeled $\ell$, then the next hole to its right is labeled $\ell+1$.
    For each $\ell\in\mathbb{Z}$, we have two Poisson clocks, with rate $p(1)$ and p(1)p(-1) respectively. All such Poisson clocks are jointly independent. The rate $p(1)$ (resp. p(1)p(-1)) clock associated to hole $\ell$ encodes the left (resp. right) jumps of this hole.
These three couplings are examples of more general exotic couplings, as defined in the introduction. We give an alternate, equivalent, description of the exotic couplings here using labels on particles and holes rather than on sites and heights.
For any ASEP configuration, we label all particles and holes by integers via the height function. More precisely, for a particle configuration with height function $h:\mathbb{Z}\to\mathbb{Z}$, if a particle is at location $x\in\mathbb{Z}$, it is given the label (h(x)+x)/2-(h(x)+x)/2; if a hole is at location $x\in\mathbb{Z}$, it is given the label (h(x)+x)/2(-h(x)+x)/2. It is straightforward to check that this labeling rule is compatible with ASEP dynamics.
@@
For any $(a,b)\in\mathbb{N}^{2}\setminus\{(0,0)\}$, we define the $(a,b)$-exotic coupling as follows. Let $\mathbb{Z}^{2}/(a,b)$ be the quotient of $\mathbb{Z}^{2}$ under addition by $(a,b)$. We associate each element $x\in\mathbb{Z}^{2}/(a,b)$ to two Poisson clocks with rate $p(1)$ and p(1)p(-1), respectively. Let $Q:\mathbb{Z}^{2}\to\mathbb{Z}^{2}/(a,b)$ denote the projection map. Any ASEP under the $(a,b)$-exotic coupling is generated by these Poisson clocks as follows. For any $\ell,k\in\mathbb{Z}$, if the particle with $\ell$ sits immediate to the left (resp. right) of the hole with label $k$, then their swap is encoded by the rate $p(1)$ (resp. p(1)p(-1)) clock associated to $Q(\ell,k)$.
@@ 7.2 AEPs under the basic coupling
In the remainder of this section, we consider general AEPs with the basic coupling. The setup is described at the beginning of Section 7. Recall from (39) that we use the notation $\mathcal{K}^{\epsilon}$ for the collection of random operators encoding the dynamics of a rescaled AEP. In this notation, the jump distribution $p$ is repressed. In this section, we will consider sequences of AEP operators $\mathcal{K}^{\epsilon}=\{\mathcal{K}^{\epsilon}_{s,t}:\operatorname{SRW}_{\epsilon}\to\operatorname{SRW}_{\epsilon},s\leq t\in\mathbb{R}\},\epsilon>0$ where the jump distributions $p_{\epsilon}$ depend on $\epsilon$, and satisfy the following:
  • For all $i$, $p_{\epsilon}(i)\to\mathbf{1}(i=1)$ as $\epsilon\to 0$. In other words, the AEPs approach TASEP as $\epsilon\to 0$.
  • There exists $K\in\mathbb{N}$ such that $p_{\epsilon}(i)=0$ for all $|i|>K$ and $|p_{\epsilon}(i)|\leq K$ for all $i\leq K$. This is a strong uniform integrability condition.
In this subsection we will prove joint convergence of AEPs to the directed landscape by appealing to Theorem 5.2. One technical claim verifying approximate monotonicity for the model is left to Section 7.3.
KPZ fixed point convergence for general AEPs in this TASEP-perturbative setting follows from the methods of [QS23] [^2] The convergence of AEPs with fixed jump distribution to the KPZ fixed point shown in [QS23] was retracted in May 2025 due to an error in the proof. Nonetheless, a modification of the proof still works in the weaker setting of Theorem 7.6 [Qua]. It will be published as an erratum to [QS23].. The convergence result applies to initial conditions which are not too far from the stationary Bernoulli product measure. We let $\nu$ be the law on $\operatorname{SRW}$ of a simple symmetric random walk path $X$ with $X(0)=0$. We let $\nu_{\epsilon}$ be the pushforward of $\nu$ under the natural map $A_{\epsilon}:\operatorname{SRW}\to\operatorname{SRW}_{\epsilon}$.
@@
  • $\limsup_{\epsilon\to 0}\|g_{\epsilon}d\nu_{\epsilon}\|_{L^{2}(\nu_{\epsilon})}<\infty$.
@@
The convergence of $\mathcal{K}^{\epsilon}_{s,t}f^{\epsilon}\stackrel{{\scriptstyle d}}{{\to}}\mathfrak{h}_{t-s}f$ is proven in [QS23], see the previous footnote. For the joint convergence, let $(f,\mathcal{K}_{s,t}f)$ be any joint subsequential limit in distribution of $(f^{\epsilon},\mathcal{K}^{\epsilon}_{s,t}f^{\epsilon})$. Let $U$ be any continuity set for the law of $f$ such that $\mathbb{P}[f\in U]>0$. Then $\mathbb{P}[f^{\epsilon}\in U]\to\mathbb{P}[f\in U]$ as $\epsilon\to 0$. Moreover, since $\mathbb{P}[f\in U]>0$, if $f^{\epsilon}_{U}$ has the law of $f^{\epsilon}$ given $f^{\epsilon}\in U$, then $f^{\epsilon}_{U}$ still satisfies the three bullets above and hence the conclusion of [QS23]. Therefore for any measurable $A\subset\operatorname{UC}$, we have $\mathbb{P}[(f,\mathcal{K}_{s,t}f)\in U\times A]=\mathbb{P}[(f,\mathfrak{h}_{t-s}f)\in U\times A]$, and so $(f,\mathcal{K}_{s,t}f)\stackrel{{\scriptstyle d}}{{=}}(f,\mathfrak{h}_{t-s}f)$ as desired. ∎
@@
Take (potentially random) initial conditions $f_{0},\ldots,f_{k}\in\operatorname{UC}$ and $f_{0}^{\epsilon},\ldots,f_{k}^{\epsilon}\in\operatorname{SRW}_{\epsilon}$ for each $\epsilon>0$, such that $f_{i}^{\epsilon}\stackrel{{\scriptstyle d}}{{\to}}f_{i}$ for each $i\in\llbracket 0,k\rrbracket$. Suppose that $f_{0}^{\epsilon}\sim\nu_{\epsilon}$ and for every $i\in\llbracket 1,k\rrbracket$, the initial conditions $f_{i}^{\epsilon}$ satisfy the last two bullet points of Theorem 7.6 and that the random variables
@@
Let $\{q_{1}<\cdots<q_{k}\}$ be the set of points where $h^{\prime}$ does not exist, so that $h^{\prime}=0$ off of $q_{1},q_{k}]$ and is constant on each interval $(q_{j},q_{j+1})$. For $i\epsilon/2\notin(q_{1},q_{k}]$ we set $v(h)_{i}=e_{i}$. Now, for each $j\in\{1,\dots,k-1\}$, let $s_{j}$ be the value of $h^{\prime}$ on $(q_{j},q_{j+1})$. For $i\epsilon/2\in(q_{j},q_{j+1}]$ we let
@@
Consider $F_{\epsilon}=A_{\epsilon}^{-1}(B_{s,n}^{\epsilon}[h])\in\operatorname{SRW}$, and let $h_{\epsilon}$ be the Radon-Nikodym derivative of its law with respect to $\nu$. Equivalently, we will show that $\|h_{\epsilon}d\nu\|_{L^{2}(\nu)}$ stays bounded as $\epsilon\to 0$. By construction, there is an integer set $Q_{\epsilon}=\{q_{1,\epsilon},\dots,q_{k,\epsilon}\}$ (up to $O(\epsilon)$-errors, this set is given by $2\epsilon^{-1}\{q_{1},\dots,q_{k}\}$ with the $q_{i}$ as above) such that conditional on $F_{\epsilon}|_{Q_{\epsilon}}$, $F_{\epsilon}$ has the same law as a simple symmetric random walk path conditioned on the values on $Q_{\epsilon}$. This follows from the fact that if we condition a parameter-$p$ Bernoulli process on $\{1,\dots,n\}$ to have exactly $k$ 11s, then under this conditioning the $k$ 11s are uniformly distributed amongst all $k$-element subsets of $\{1,\dots,n\}$; in particular, the parameter $p$ plays no role.
@@
The final two assumptions in Theorem 5.2 require a monotonicity claim for general AEPs.
For this next two lemmas, we consider AEPs where the jump distribution $p$ satisfies that $p(1)>1/2$, $p(i)=0$ for all $|i|>K$, and $|p(i)|\leq K$ for all $i\leq K$, for some $K\in\mathbb{N}$. Take two AEPs with the same such $p$, with height functions $(h^{-}_{t})_{t\geq 0}$ and $(h^{+}_{t})_{t\geq 0}$, started at time 0 in the basic coupling. Assume that $h^{-}_{0}\leq h^{+}_{0}$. For any $t,w>0$, we denote
@@
For any $m,t,w>0$, we have $\mathbb{P}[Y_{t,w}>m]<C(t+w)\exp(-cm)+C\exp(-ct)$, where $C,c>0$ are constants depending only on $K$.
@@
Suppose that $\eta^{1}_{t},\eta^{2}_{t}$ are two AEPs with the basic coupling started from initial conditions $\eta^{1}_{0},\eta^{1}_{0}$ that agree off of a set $\{-n,\dots,n\}$. Then there exists $C_{1},C_{2}>0$ (depending only on $K$) such that for all $t>0$,
@@
Let $M_{t}=\sup\{m\in\mathbb{N}:(\eta^{1}_{t}(-m),\eta^{1}_{t}(m))\neq(\eta^{2}_{t}(-m),\eta^{2}_{t}(m))\}$. Then the process $(M_{t},\eta^{1}_{t},\eta^{2}_{t})$ is jointly Markov, and since the jump distribution $p$ has finite support, the process $M_{t}$ is stochastically dominated by $N_{t}+n$, where $N_{t},t\geq 0$ is a Poisson counting process with rate $K^{2}$. The bound in the lemma then follows from a standard tail estimate on Poisson random variables. ∎
@@
    $\mathcal{K}_{t_{*},t}(B_{t_{*},1}+h_{m})\geq\tilde{\mathcal{K}}_{0,t}f_{i}\qquad$ $\text{for each }m\in\mathbb{N},t\in T_{*},t>t_{*}.$
@@
where $M_{\epsilon},\epsilon>0$ is a tight collection of random variables and $N_{\epsilon}\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}0$ with $\epsilon$. Here the convergence of $N_{\epsilon}$ to 0 uses Lemma 7.10 and the tightness of $M_{\epsilon}$ uses Lemma 7.9. The inequality (43) follows by taking $\epsilon\to 0$.
@@
is tight. If we take any subsequential limit $S$, jointly with (42), then by applying Lemmas 7.9 and 7.10 as before we would have that $\mathcal{K}_{r,t}g\leq\mathcal{K}_{s,t}f+S$. This would imply part $2$ if we could show that $S=\sup(g-\mathcal{K}_{s,r}f)$. However, this is not obvious and so we work around this difficulty with a truncation argument.
For each $m\in\mathbb{N}$, define $w_{m}\in\mathscr{H}$ by letting $w_{m}(x)=-1/m$ when $x\in[-m,m]$, and $w_{m}(x)=-m$ when $x\in[-m-1,m+1]^{c}$, and letting $w_{m}$ be linear on [m1,m][-m-1,-m] and $[m,m+1]$. Then similarly, the collection of random variables
@@
with probability $\to 1$ as $\epsilon\to 0$. Then by using Lemmas 7.9 and 7.10 as before we have that
@@
For part $3$, we let $R^{\epsilon}=\sup(f_{i}^{\epsilon}-f_{0}^{\epsilon})$ and note that $(R^{\epsilon})_{\epsilon>0}$ is again tight, as assumed in Theorem 7.7. Then using Lemmas 7.9 and 7.10 as in previous parts,
@@
where $E_{\epsilon}\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}0$ as $\epsilon\to 0$. Note that $\mathcal{K}_{0,t_{*}}^{\epsilon}f_{0}^{\epsilon}(0)$ is also tight according to Theorem 7.6. Now let $R$ be a joint subsequential limit of $R^{\epsilon}+\mathcal{K}_{0,t_{*}}^{\epsilon}f_{0}^{\epsilon}(0)$ taken with the other random variables. The above display together with the monotonicity argument from previous parts (again using Lemmas 7.9 and 7.10) implies that if $h\in\mathscr{H}$ is any function with $h(x)\geq R+1$ for all large enough $|x|$, and $B_{t_{*},1}(x)+h(x)>\tilde{\mathcal{K}}_{0,t_{*}}f_{i}(x)$ for all $x$, then
@@
for all $\in T_{*}, t>t$$t_{*}$. Finally, by choosing any decreasing sequence $\{h_{m}\}_{m\in\mathbb{N}}$ satisfying $B_{t_{*},1}+h_{m}\downarrow\tilde{\mathcal{K}}_{0,t_{*}}f_{i}$ yields the conclusion. (Here the existence of such a sequence is guaranteed by the fact that both $B_{t_{*},1}$ and $\mathcal{K}_{0,t_{*}}f_{i}$ are continuous, due to the fact that $\mathcal{K}_{0,t_{*}}f_{i}\stackrel{{\scriptstyle d}}{{=}}\mathfrak{h}_{t_{*}}f_{i}$ by Theorem 7.6, and that $B_{t_{*},1}$ is a Brownian motion.) ∎